The Morris Corporation has $750,000 of debt outstanding, and it pays an interest rate of 8% annually. Morris's annual sales are $3.75 million, its average tax rate is 35%, and its net profit margin on sales is 5%. If the company does not maintain a TIE ratio of at least 4 to 1, its bank will refuse to renew the loan and bankruptcy will result. What is Morris's TIE ratio? Round intermediate calculations to two decimal places. Round your answer to two decimal places.
Answer #1

interest amount = 750000*.08 = 60000

net profit = 5% of 3.75 million = 187500

net profit + tax = profit before tax = 187500/.65 = 288,461.54

profit before tax + interest = earning before interest and tax = 288461.54+60000 = 348461.54

TIE ratio = EBIT/Interest = 348461.54/60000 = 5.81

bank will not refuse to renew

Thanks

Let me know if you have any doubt.